OpenStudy (anonymous):

IM GIVING MEDALS TO SOMEONE WHO ANSWERS THIS! Will someone write the function for when the vertex of a parabola is at (2,3) and passes through the points (1,2) and (1,4) thanks(:

OpenStudy (anonymous):

@phi @hero @mathstudent55

OpenStudy (hlambach):

You have three points, thus you have three formulas. \[3=a2^{2}+b(2)+c \] \[2=a1^{2}+b(1)+c\] \[4=a1^{2}+b(1)+c\]

OpenStudy (anonymous):

@hlambach It just shows a graph and I need to know the function of it to compare it to my other functions. It's supposed to equal h(x). I probably should have included that

OpenStudy (anonymous):

This is terrible quality but here is a pic

OpenStudy (anonymous):

Please help ): @hlambach

OpenStudy (anonymous):

@igreen

OpenStudy (hlambach):

Sorry, my internet is really slow right now.

OpenStudy (hlambach):

So it gives you the points?

OpenStudy (hlambach):

You can right it in vertex form and then FOIL. Vertex form is \[y=a(x-h)^{2}+k\] (-h,k) is your vertex. The vertex in the graph is (3,2) Substitute it in.

OpenStudy (anonymous):

@hlambach thank you omg hang on

OpenStudy (hlambach):

*write

OpenStudy (hlambach):

haha, no proble.

OpenStudy (anonymous):

@hlambach so h(x)=a(x-3)^2=2?

OpenStudy (hlambach):

The formula is a(x-3)^2=2. FOIL that. So multiply it out.

OpenStudy (anonymous):

Take the = 2 off the end I have no idea why I put that

OpenStudy (hlambach):

Thanks for the medal. :)

OpenStudy (hlambach):

I'm so sorry, wait. Before you multiply it out. You need to find a. To find a, use any point on the graph (that isn't the vertex) and put it in for x and y.

OpenStudy (anonymous):

@hlambach it's okay haha thank you so much I've been stuck on this forever

OpenStudy (hlambach):

So solve for a in this equation (1,2) 2=a(1-3)^2+2. (sorry, I forgot to add the +2 on the previous equation.)

OpenStudy (hlambach):

Once you get that, change the equation back to y=(whatever you get for a)(x-3)^2+2 and multiply it all out.

OpenStudy (hlambach):

Once that's done just change the y to h(x). (It's the same thing.

OpenStudy (hlambach):

Get it? @Kaitlynbrooks33

OpenStudy (anonymous):

@hlambach Yes thank you(: is there anyway to give multiple medals lol

OpenStudy (hlambach):

haha, I wish there were. Glad I could help! Make sure to close the question!